A neutron is a spherical object that has a mass of 0.000000000000000000000000001675 kilograms and a radius of 0.8 femtometers. The volume of a sphere is (4/3)(π)r3. What is the density of a proton?
1
Expert's answer
2020-06-15T14:38:13-0400

Density of a proton is defined as'

"Density (\rho)=\dfrac{Mass(M)}{Volume(V)}" and'


"Volume(V)=(\dfrac{4}{3})\pi r^{3}"

We are given radius (r) =0.8fm

Converting into metres.

"1fm=1.0*10^{-15}m"

"0.8fm=??"

"=8*10^{-16}m"


Therefore,

"V=1.33*3.14*(8*10^{-16})^{3}"

"=2.138*10^{-45}m^3"

So,

"\rho=\dfrac{1.675*10^{27}}{2.138*10^{-45}}"


"=7.834*10^{71}kg\/m^3"